Let * be the binary operation defined on Q. Find which of the following binary operations are commutative   
a * b = a + ab ∀ a, b ∈ Q  

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

a * b = a + ab

b * a = b + ab
clearly it can be seen that,
a + ab ≠ b + ab
Hence , * is not commutative.

Thanks